Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Problem Solving — algebra and arithmetic
Expert replies
by andressa » Wed Feb 02, 2011 6:58 am
Hi, can someone explain to me how to work on this two problems? I tried to work on them but I cant find the right answer.

There are 3 red balls, and 2 black balls in a jar. What is the probability that John gets at least one red ball when he picks two balls at random out of the jar at the same time?
a) 6/25
b) 19/25
c) 1/10
d) 9/10
e) 19/20

In a bowl of marbles, 8 are yellow, 6 are blue, and 4 are black.. If John picks 2 balls out of the bowl at random and at the same time, what is the probability that at least one of the marbles will be yellow?
a)5/17
b)12/17
c)25/81
d)56/81
e)4/9
Join the discussion
Source: — Quantitative Reasoning |

by Geva@EconomistGMAT » Wed Feb 02, 2011 7:15 am
andressa wrote:Hi, can someone explain to me how to work on this two problems? I tried to work on them but I cant find the right answer.

There are 3 red balls, and 2 black balls in a jar. What is the probability that John gets at least one red ball when he picks two balls at random out of the jar at the same time?
a) 6/25
b) 19/25
c) 1/10
d) 9/10
e) 19/20

P(at at least one red) has too many scenarios: red only on first, red only on second, red on both.
On the other end, the only forbidden scenario is "no red" = two blacks. Go for 1-p(forbidden scenario) = 1-p(two blacks) = 1-2/5*1/4 = 1-1/10 = 9/10. Answer is D.

Generally speaking, at least questions lend themselves to the 1-p(forbidden scenario) approach.


In a bowl of marbles, 8 are yellow, 6 are blue, and 4 are black.. If John picks 2 balls out of the bowl at random and at the same time, what is the probability that at least one of the marbles will be yellow?
a)5/17
b)12/17
c)25/81
d)56/81
e)4/9
Same as above: P(at least one yellow) = 1-p(no yellow)
find forbidden p(not yellow) =this is the same as P(no yellow on first) * p(no yellow on second) = 10/18 * 9/17 = 10/2 * 1/17 = 5/17
So p(at least one yellow) = 1-p(no yellow) = 1-5/17 = 12/17. Answer is B.
Geva
Senior Instructor
Master GMAT
1-888-780-GMAT
https://www.mastergmat.com
Join the discussion

by andressa » Wed Feb 02, 2011 9:43 am
Thanks a bunch....I was getting confused on the calculation wise!! Argh....sometimes the frustration gets on the way.
Join the discussion